OpenStudy (anonymous):

If you have 1.5g ordered, and you have 15mg/mL on hand. How many mL's do you need to give th ordered amount? (work shown is needed)

OpenStudy (anonymous):

1g = 1000mg so 1.5g = 1.5 * 1000mg = 1500mg Now you have 15 mg onhand, so the remaining amount that you need to get is 1500 - 15 = 1485mg